# Markov Chains Approximate Message Passing

### v1

@@ Abstract
Markov chain Monte Carlo algorithms have long been observed to obtain near-optimal performance in various Bayesian inference settings. However, developing a supporting theory that make these studies rigorous has proved challenging.
@@ 1 Introduction
For certain inference problems, one can use state evolution to design AMP algorithms to construct Bayes-optimal estimators. Similarly, we a priori know Glauber achieves the same guarantees in exponential time, since by then the dynamics would be able to produce Gibbs samples. Motivated by this, we refine our initial question to ask about connecting the performance of these two classes of algorithms if Glauber is only run for polynomial time.
@@
The precision of our results require model-specific calculations that we did not attempt to generalize beyond the spiked Wigner setting. However, our high-level proof strategy is fairly generic and we believe it should be applicable to general spiked matrix inference problems.
@@
For the spiked Wigner problem, they show that for sufficiently large (constant) $\lambda$ and $\beta<\frac{1}{4}$, the Glauber dynamics run on $\mu_{\beta\bm{M}}$ for a large polynomial time from an arbitrary initialization reaches a vector that is non-trivially correlated with the spike. To prove this result, they combined the properties of locally stationary distributions with an analysis of a different Markov chain known as Restricted Gaussian Dynamics (RGD), which we define below. [^1] The precise definition is not terribly important, so readers seeking intuition can skip the definition without much concern.
@@ 1.1 Main results
The following theorem establishes that the short-term behavior of the correlation of Glauber iterates are approximate fixed points of an AMP-like fixed point equation in the entire high SNR regime $\lambda>1$. Furthermore, it establishes a sharp phase transition at $\beta=\frac{1}{\lambda}$ for the success of Glauber dynamics at weak recovery of the spike.
@@ 2 Technical Overview
The technical overview is organized as follows. We will first define RGD and describe the high-level reduction of why one should expect RGD to boil down to a one-dimensional recursion on correlation. Then, in Section 2.1, we show how spin glass machinery can be leveraged to explicitly understand what this one-dimensional recursion is. Once we have written down the explicit recursion, we can readily connect it to AMP; this is carried out in Section 2.2. Finally, we describe some qualitative properties of this recursion, such as the nature of its fixed points, in Section 2.3.
First off, observe that due to the rotational symmetry of the noise matrix and prior, we may assume without loss of generality that the spike is the all-ones vectors $\bm{1}$. In this case, the scaled posterior (2) of the spiked Wigner model becomes
@@
We start by explaining where RGD arises from, and the connection between the posterior and the SK model. It arises from the Hubbard–Stratonovich transform [Str57, Hub59] (also see [LMRW24, Theorem 3.12]). Let $\bm{\sigma}\sim\mu_{\beta W+\frac{\beta\lambda}{N}\bm{1}\bm{1}^{\top}}$, and $\bm{g}\sim\mathcal{N}\left(0,1\right)$. Set $\bm{z}=\frac{\beta\lambda}{N}\langle\bm{\sigma},\bm{1}\rangle+\sqrt{\frac{\beta\lambda}{N}}\bm{g}$. Then, the posterior distribution on $\bm{\sigma}$ conditioned on $\bm{z}$ is simply
$\mathbf{Pr}\left[\sigma\mid\bm{z}\right]$ $\propto\mu_{\beta\bm{M}}(\sigma)\cdot\exp\left(-\frac{N}{2\beta\lambda}\left(\bm{z}-\frac{\beta\lambda}{N}\langle\sigma,\bm{1}\rangle\right)\right)$
@@
$\mu_{\beta W+\frac{\beta\lambda}{N}\bm{1}\bm{1}^{\top}}=\bm{\mathrm{E}}_{\bm{z}}\mu_{\beta W,\bm{z}\bm{1}}\,.$
@@
This motivates us to consider the related $P^{\mathrm{RGD}}$ Markov chain on [1,1][-1,1], which tracks the (normalized) correlation of the $N$-dimensional RGD iterate with the spike. In this Markov chain, if starting at $y\in\mathbb{R}$, we draw $\bm{g}\sim\mathcal{N}\left(0,1\right)$, set $\bm{z}=\beta\lambda y+\sqrt{\frac{\beta\lambda}{N}}\bm{g}$, draw $\bm{\sigma}\sim\mu_{\beta\bm{W},\bm{h}\bm{1}}$, and finally move to $\bm{z}^{\prime}=\frac{1}{N}\langle\bm{\sigma},\bm{1}\rangle$. The stationary distribution of this Markov chain is just the (normalized) marginal of $\mu_{\beta\bm{M}}$ along the all-ones direction.
@@
2.1 The mean recursion and fixed points structure
@@
Define the mean magnetization $\bm{m}_{h}=\frac{1}{N}\bm{\mathrm{E}}_{\bm{\sigma}\sim\mu_{\beta W,h}}\langle\bm{\sigma},\bm{1}\rangle$, and for $\bm{\sigma}\sim\mu_{\beta W,h}$, let $\widetilde{\bm{g}}_{h}$ be the random variable $\sqrt{N}\cdot\left(\frac{1}{N}\langle\bm{\sigma},\bm{1}\rangle-\bm{m}_{h}\right)$. Under this notation, we can write down the following description of $P^{\mathrm{RGD}}$: from $z\in\mathbb{R}$,
@@
To simplify the recursion further, we need concentration of $\bm{z}^{\prime}$ around $\bm{m}_{h}$, as well as an explicit formula for $\bm{m}_{h}$. The key observation is that such properties are tractable to prove if the SK model $\mu_{\beta W,h}$ is in the high temperature regime. We first need to introduce some standard spin glass quantities.
@@
We show that if (AT) is satisfied, the correlation $\langle\bm{\sigma},\bm{1}\rangle$ of a Gibbs sample $\bm{\sigma}\sim\mu_{\beta W,h}$ concentrates around the mean correlation. In other words, we have $\bm{z}^{\prime}\approx\bm{m}_{\bm{h}}$ with high probability. Hence, we conclude that a single step of $P^{\mathrm{RGD}}$ can be approximated by the map $z\mapsto\bm{m}_{\beta\lambda z}$.
@@ 2.3 Qualitative behavior of the fixed point equations
Convergence is nontrivial to guarantee, and indeed the proof of convergence for Bayes AMP does not immediately generalize to the mismatched setting where $\beta\neq\lambda$. To this end, in Proposition 7.1 we establish that so long as $(\beta,\beta\lambda z)$ satisfy (AT), the map $z\mapsto\bm{m}_{\beta\lambda z}$ is strictly increasing in $z$. For any $\beta<1$, (AT) is automatically satisfied, and in this setting we are able to fully characterize convergence for the recursion starting from any $z$.
@@
    If $\beta>1$, the formula for the mean correlation $\bm{m}_{\beta\lambda z}$ is only valid if $z$ is sufficiently large. Indeed, if $z$ is close to 0, (AT) is violated, and the corresponding SK model is known to be in the low-temperature replica-symmetry breaking (RSB) regime. Hence, the recursion is only valid for large $z$, and in this region there are at most two stable fixed points $\pm\mathsf{OPT}_{\beta}$.
@@
    First, that for any $\beta$, the RGD Markov chain reaches the stable $\mathsf{OPT}_{\beta}$ fixed point from a sufficiently warm start which puts $z_{0}$ in the high-temperature, replica-symmetric (RS) regime. That is, if the initialization $z_{0}$ of $P^{\mathrm{RGD}}$ is such that $(\beta,\beta\lambda z_{0})$ satisfies (AT), then the correlation shoots to the stable fixed point $\mathsf{OPT}_{\beta}$ in a small constant number of steps of RGD. This is essentially immediate from classical analyses of fixed point recursions, see Lemma 5.4 for details.
@@
    Second, that for $\beta<1$, the RGD Markov chain reaches the stable $\mathsf{OPT}_{\beta}$ fixed point from an arbitrary initialization. In the case that $\beta<\frac{1}{\lambda}$, 0 is the only stable fixed point, so for now we assume that $\beta\in(\frac{1}{\lambda},1)$. Since the previous point implies that the correlation goes to $\mathsf{OPT}_{\beta}$ once it becomes sufficiently large, the interesting part here is showing that the Markov chain successfully escapes the unstable fixed point at 0. We establish this in Lemma 5.6, where we show that the noise $\bm{g}$ and $\widetilde{\bm{g}}$, which are typically $O\left(\frac{1}{\sqrt{N}}\right)$, are anticoncentrated enough to cause the correlation to escape 0.
@@ 3 Conjectures and open questions
We next turn to open questions regarding the spiked Wigner problem in particular. One fundamental barrier to applying our techniques to understand Glauber for $\beta>1$ is that the transfer of local stationarity (Lemma 4.8) does not work in this setting, even if we assume mixing guarantees in the entire conjecture high-temperature region (AT). In particular, it is unclear how to prove a version of Lemma 4.8 where we only assume that the component measure satisfies an MLSI with high probability, under the assumption that the initialization from which the locally stationary distribution is obtained lives in this high-temperature region.
@@
Assuming mixing for SK everywhere where $(\beta,h)$ satisfy (AT), can we extend our analysis to characterize the recovery performance for annealed Glauber dynamics up to $\beta\in\left(1,\beta_{\lambda}\right)$ for some $\beta_{\lambda}>\lambda$?
@@
One path towards achieving such a guarantee is to establish exponential contraction of the local stationarity parameter, perhaps by establishing a weak Poincaré inequality [HMRW24] for symmetric test functions. Indeed, we believe that our techniques to analyze RGD might help in showing that annealed Glauber dynamics succeeds at sampling from the posterior (and potentially even from lower-temperature models, up to the aforementioned threshold $\beta_{\lambda}>\lambda$). Indeed, while the “conservation of variance” property used to show sampling guarantees for annealed Glauber dynamics is typically reliant on proving certain covariance bounds (e.g. [HMRW24, Lemma 6.8]) which is often technically challenging, perhaps this can be sidestepped if one instead directly tried to establish a weak Poincaré inequality for the (one-dimensional!) RGD Markov chain.
@@ 4.2 Markov chain basics
We also recall the standard definitions of functional inequalities which imply rapid mixing. For any functions $f,g:\Omega\to\mathbb{R}$, define the Dirichlet form $\mathcal{E}(f,g)\coloneqq\bm{\mathrm{E}}_{{\bm{x}}\sim\pi}\bm{\mathrm{E}}_{\bm{y}\sim_{P}{\bm{x}}}(f({\bm{x}})-f(\bm{y}))(g({\bm{x}})-g(\bm{y}))$.
@@ 4.3 Locally stationary distributions
Let $P$ be a time-reversible ergodic Markov chain on a finite state space $\Omega$, with associated stationary distribution $\pi$. A probability measure $\nu$ on $\Omega$ is said to be $\varepsilon$-locally stationary with respect to $P$ if
@@
Let $\nu_{0}$ be an arbitrary distribution, $\nu_{t}=P_{t}\nu_{0}$, and set $f_{t}=\frac{\mathop{}\!\mathrm{d}\nu_{t}}{\pi}$. Also set $f=\bm{\mathrm{E}}_{\bm{t}\sim[0,T]}f_{\bm{t}}$. Then,
@@
In particular, $\nu\coloneqq\bm{\mathrm{E}}_{\bm{t}\sim[0,T]}\nu_{\bm{t}}$ is $\frac{\log(1/\pi_{\mathrm{min}})}{T}$-locally stationary with respect to $\pi$.
@@ 4.4 Approximate Message Passing and State Evolution
State evolution [BM11] predicts the empirical distribution of the coordinates of $x_{t}$. We record the special case relevant to our setting: the correlation of the AMP iterate with the planted signal $\bm{x}$} can be predicted by a scalar recursion.
@@ 4.5 The Sherrington–Kirkpatrick model
Significant progress has been made towards establishing that for the SK model, $(\beta,h)$ satisfies (AT) if and only if the corresponding model is replica-symmetric. On the negative side, it was shown by Toninelli [Ton02] that if $(\beta,h)$ does not satisfy (AT), the corresponding model is replica-symmetric breaking. On the positive side, it was shown by Talagrand [Tal11, Theorem 13.7.1] that exponential overlap concentration holds for a very large subregion of the area above the AT line, which is conjectured to coincide with the entirety of the region above the AT line. Making this conjecture rigorous requires proving an explicit numerical condition. We refrain from elaborating as the details are not directly relevant to our work, and instead refer the interested reader to Talagrand’s text, as well as the discussion in [JT17, p.8]. We remark that it is known that Talagrand’s subregion contains the entire high-temperature regime $\beta<1$, see e.g. [BXY23, p.3] and [Che21].

### v2

@@ Abstract
Markov chain Monte Carlo algorithms have long been observed to obtain near-optimal performance in various Bayesian inference settings. However, developing a supporting theory that makes these studies rigorous has proved challenging.
@@ 1 Introduction
For certain inference problems, one can use state evolution to design AMP algorithms to construct Bayes-optimal estimators. Similarly, we a priori know that Glauber achieves the same guarantees in exponential time, since by then the dynamics would be able to produce Gibbs samples. Motivated by this, we refine our initial question to ask about connecting the performance of these two classes of algorithms if Glauber is only run for polynomial time.
@@
The precision of our results requires model-specific calculations that we did not attempt to generalize beyond the spiked Wigner setting. However, our high-level proof strategy is fairly generic and we believe it should be applicable to general spiked matrix inference problems.
@@
For the spiked Wigner problem, they show that for sufficiently large (constant) $\lambda$ and $\beta<\frac{1}{4}$, the Glauber dynamics run on $\mu_{\beta\bm{M}}$ for a large polynomial time from an arbitrary initialization reaches a vector that is nontrivially correlated with the spike. To prove this result, they combined the properties of locally stationary distributions with an analysis of a different Markov chain known as Restricted Gaussian Dynamics (RGD), which we define below. [^1] The precise definition is not terribly important, so readers seeking intuition can skip the definition without much concern.
@@ 1.1 Main results
The following theorem establishes that the short-term behavior of the correlation of Glauber iterates is an approximate fixed point of an AMP-like fixed point equation in the entire high SNR regime $\lambda>1$. Furthermore, it establishes a sharp phase transition at $\beta=\frac{1}{\lambda}$ for the success of Glauber dynamics at weak recovery of the spike.
@@ 2 Technical Overview
The technical overview is organized as follows. We will first define RGD and describe the high-level reduction showing why one should expect RGD to boil down to a one-dimensional recursion on correlation. Then, in Section 2.1, we show how spin glass machinery can be leveraged to explicitly understand what this one-dimensional recursion is. Once we have written down the explicit recursion, we can readily connect it to AMP; this is carried out in Section 2.2. Finally, we describe some qualitative properties of this recursion, such as the nature of its fixed points, in Section 2.3.
First off, observe that due to the rotational symmetry of the noise matrix and prior, we may assume without loss of generality that the spike is the all-ones vector $\bm{1}$. In this case, the scaled posterior (2) of the spiked Wigner model becomes
@@
We start by explaining where RGD arises from, and the connection between the posterior and the SK model. It arises from the Hubbard–Stratonovich transform [Str57, Hub59] (also see [LMRW24, Theorem 3.12]). Let $\bm{\sigma}\sim\mu_{\beta\bm{W}+\frac{\beta\lambda}{N}\bm{1}\bm{1}^{\top}}$, and $\bm{g}\sim\mathcal{N}\left(0,1\right)$. Set $\bm{z}=\frac{\beta\lambda}{N}\langle\bm{\sigma},\bm{1}\rangle+\sqrt{\frac{\beta\lambda}{N}}\bm{g}$. Then, the posterior distribution on $\bm{\sigma}$ conditioned on $\bm{z}$ is simply
$\mathbf{Pr}\left[\sigma\mid\bm{z}\right]$ $\propto\mu_{\beta\bm{M}}(\sigma)\cdot\exp\left(-\frac{N}{2\beta\lambda}\left(\bm{z}-\frac{\beta\lambda}{N}\langle\sigma,\bm{1}\rangle\right)^{2}\right)$
@@
$\mu_{\beta\bm{W}+\frac{\beta\lambda}{N}\bm{1}\bm{1}^{\top}}=\bm{\mathrm{E}}_{\bm{z}}\mu_{\beta\bm{W},\bm{z}\bm{1}}\,.$
@@
This motivates us to consider the related $P^{\mathrm{RGD}}$ Markov chain on [1,1][-1,1], which tracks the (normalized) correlation of the $N$-dimensional RGD iterate with the spike. In this Markov chain, if starting at $y\in\mathbb{R}$, we draw $\bm{g}\sim\mathcal{N}\left(0,1\right)$, set $\bm{z}=\beta\lambda y+\sqrt{\frac{\beta\lambda}{N}}\bm{g}$, draw $\bm{\sigma}\sim\mu_{\beta\bm{W},\bm{z}\bm{1}}$, and finally move to $\bm{z}^{\prime}=\frac{1}{N}\langle\bm{\sigma},\bm{1}\rangle$. The stationary distribution of this Markov chain is just the (normalized) marginal of $\mu_{\beta\bm{M}}$ along the all-ones direction.
@@
2.1 The mean recursion and fixed point structure
@@
Define the mean magnetization $\bm{m}_{h}=\frac{1}{N}\bm{\mathrm{E}}_{\bm{\sigma}\sim\mu_{\beta\bm{W},h}}\langle\bm{\sigma},\bm{1}\rangle$, and for $\bm{\sigma}\sim\mu_{\beta\bm{W},h}$, let $\widetilde{\bm{g}}_{h}$ be the random variable $\sqrt{N}\cdot\left(\frac{1}{N}\langle\bm{\sigma},\bm{1}\rangle-\bm{m}_{h}\right)$. Under this notation, we can write down the following description of $P^{\mathrm{RGD}}$: from $z\in\mathbb{R}$,
@@
To simplify the recursion further, we need concentration of $\bm{z}^{\prime}$ around $\bm{m}_{h}$, as well as an explicit formula for $\bm{m}_{h}$. The key observation is that such properties are tractable to prove if the SK model $\mu_{\beta\bm{W},h}$ is in the high temperature regime. We first need to introduce some standard spin glass quantities.
@@
We show that if (AT) is satisfied, the correlation $\langle\bm{\sigma},\bm{1}\rangle$ of a Gibbs sample $\bm{\sigma}\sim\mu_{\beta\bm{W},h}$ concentrates around the mean correlation. In other words, we have $\bm{z}^{\prime}\approx\bm{m}_{\bm{h}}$ with high probability. Hence, we conclude that a single step of $P^{\mathrm{RGD}}$ can be approximated by the map $z\mapsto\bm{m}_{\beta\lambda z}$.
@@ 2.3 Qualitative behavior of the fixed point equations
Convergence is nontrivial to guarantee, and indeed the proof of convergence for Bayes AMP does not immediately generalize to the mismatched setting where $\beta\neq\lambda$. To this end, in Proposition 7.1 we establish that so long as $(\beta,\beta\lambda z)$ satisfies (AT), the map $z\mapsto\bm{m}_{\beta\lambda z}$ is strictly increasing in $z$. For any $\beta<1$, (AT) is automatically satisfied, and in this setting we are able to fully characterize convergence for the recursion starting from any $z$.
@@
    If $\beta>1$, the formula for the mean correlation $\bm{m}_{\beta\lambda z}$ is only valid if $z$ is sufficiently large. Indeed, if $z$ is close to 0, (AT) is violated, and the corresponding SK model is known to be in the low-temperature replica-symmetry breaking (RSB) regime. Hence, the recursion is only valid for large $z$, and in this region there are at most two stable fixed points $\pm\mathsf{OPT}_{\beta,\lambda}$.
@@
    First, that for any $\beta$, the RGD Markov chain reaches the stable $\mathsf{OPT}_{\beta,\lambda}$ fixed point from a sufficiently warm start which puts $z_{0}$ in the high-temperature, replica-symmetric (RS) regime. That is, if the initialization $z_{0}$ of $P^{\mathrm{RGD}}$ is such that $(\beta,\beta\lambda z_{0})$ satisfies (AT), then the correlation shoots to the stable fixed point $\mathsf{OPT}_{\beta,\lambda}$ in a small constant number of steps of RGD. This is essentially immediate from classical analyses of fixed point recursions; see Lemma 5.4 for details.
@@
    Second, that for $\beta<1$, the RGD Markov chain reaches the stable $\mathsf{OPT}_{\beta,\lambda}$ fixed point from an arbitrary initialization. In the case that $\beta<\frac{1}{\lambda}$, 0 is the only stable fixed point, so for now we assume that $\beta\in(\frac{1}{\lambda},1)$. Since the previous point implies that the correlation goes to $\mathsf{OPT}_{\beta,\lambda}$ once it becomes sufficiently large, the interesting part here is showing that the Markov chain successfully escapes the unstable fixed point at 0. We establish this in Lemma 5.6, where we show that the noise $\bm{g}$ and $\widetilde{\bm{g}}$, which are typically $O\left(\frac{1}{\sqrt{N}}\right)$, are anticoncentrated enough to cause the correlation to escape 0.
@@ 3 Conjectures and open questions
We next turn to open questions regarding the spiked Wigner problem in particular. One fundamental barrier to applying our techniques to understand Glauber for $\beta>1$ is that the transfer of local stationarity (Lemma 4.8) does not work in this setting, even if we assume mixing guarantees in the entire conjectured high-temperature region (AT). In particular, it is unclear how to prove a version of Lemma 4.8 where we only assume that the component measure satisfies an MLSI with high probability, under the assumption that the initialization from which the locally stationary distribution is obtained lives in this high-temperature region.
@@
Assuming SK mixes fast for all $(\beta,h)$ satisfying (AT), can we extend our analysis to characterize the recovery performance for annealed Glauber dynamics up to $\beta\in\left(1,\beta_{\lambda}\right)$ for some $\beta_{\lambda}>\lambda$?
@@
One path towards achieving such a guarantee is to establish exponential contraction of the local stationarity parameter, perhaps by establishing a weak Poincaré inequality [HMRW24] for symmetric test functions. Indeed, we believe that our techniques to analyze RGD might help in showing that annealed Glauber dynamics succeeds at sampling from the posterior (and potentially even from lower-temperature models, up to the aforementioned threshold $\beta_{\lambda}>\lambda$). Furthermore, while the “conservation of variance” property used to show sampling guarantees for annealed Glauber dynamics is typically reliant on proving certain covariance bounds (e.g. [HMRW24, Lemma 6.8]) which are often technically challenging, perhaps this can be sidestepped if one instead directly tried to establish a weak Poincaré inequality for the (one-dimensional!) RGD Markov chain.
@@ 4.2 Markov chain basics
We also recall the standard definitions of functional inequalities which imply rapid mixing. For a reversible Markov chain $P$ on state space $\Omega$ with stationary distribution $\pi$, and any functions $f,g:\Omega\to\mathbb{R}$, define the Dirichlet form $\mathcal{E}(f,g)\coloneqq\bm{\mathrm{E}}_{{\bm{x}}\sim\pi}\bm{\mathrm{E}}_{\bm{y}\sim_{P}{\bm{x}}}(f({\bm{x}})-f(\bm{y}))(g({\bm{x}})-g(\bm{y}))$.
@@ 4.3 Locally stationary distributions
[…]
@@
Let $\nu_{0}$ be an arbitrary distribution, $\nu_{t}=P^{t}\nu_{0}$, and set $f_{t}=\frac{\mathop{}\!\mathrm{d}\nu_{t}}{\mathop{}\!\mathrm{d}\pi}$. Also set $f=\bm{\mathrm{E}}_{\bm{t}\sim[0,T]}f_{\bm{t}}$, and $\nu\coloneqq\bm{\mathrm{E}}_{\bm{t}\sim[0,T]}\nu_{\bm{t}}$. Then,
@@
In particular, $\nu$ is $\frac{\log(1/\pi_{\mathrm{min}})}{T}$-locally stationary with respect to $\pi$.
@@ 4.4 Approximate Message Passing and State Evolution
State evolution [BM11] predicts the empirical distribution of the coordinates of $x^{t}$. We record the special case relevant to our setting: the correlation of the AMP iterate with the planted signal $\bm{x}$} can be predicted by a scalar recursion.
@@ 4.5 The Sherrington–Kirkpatrick model
Significant progress has been made towards establishing that for the SK model, $(\beta,h)$ satisfies (AT) if and only if the corresponding model is replica-symmetric. On the negative side, it was shown by Toninelli [Ton02] that if $(\beta,h)$ does not satisfy (AT), the corresponding model exhibits replica symmetry breaking. On the positive side, it was shown by Talagrand [Tal11, Theorem 13.7.1] that exponential overlap concentration holds for a very large subregion of the area above the AT line, which is conjectured to coincide with the entirety of the region above the AT line. Making this conjecture rigorous requires proving an explicit numerical condition. We refrain from elaborating as the details are not directly relevant to our work, and instead refer the interested reader to Talagrand’s text, as well as the discussion in [JT17, p.8]. We remark that it is known that Talagrand’s subregion contains the entire high-temperature regime $\beta<1$, see e.g. [BXY23, p.3] and [Che21].